### v2

# Weijia Shi1Jaechan Lee1Yangsibo Huang2 Sadhika Malladi2Jieyu Zhao3Ari Holtzman4Daogao Liu1 Luke Zettlemoyer1Noah A. Smith1Chiyuan Zhang5

Abstract
Language models (LMs) are trained on vast amounts of text data, which may include private and copyrighted content, and data owners may request the removal of their data from a trained model due to privacy or copyright concerns. However, exactly unlearning only these datapoints (i.e., retraining with the data removed) is intractable in modern-day models, leading to the development of many approximate unlearning algorithms. Evaluation of the efficacy of these algorithms has traditionally been narrow in scope, failing to precisely quantify the success and practicality of the algorithm from the perspectives of both the model deployers and the data owners. We address this issue by proposing MUSE, a comprehensive machine unlearning evaluation benchmark that enumerates six diverse desirable properties for unlearned models: (1) no verbatim memorization, (2) no knowledge memorization, (3) no privacy leakage, (4) utility preservation on data not intended for removal, (5) scalability with respect to the size of removal requests, and (6) sustainability over sequential unlearning requests. Using these criteria, we benchmark how effectively eight popular unlearning algorithms on 7B-parameter LMs can unlearn Harry Potter books and news articles. Our results demonstrate that most algorithms can prevent verbatim memorization and knowledge memorization to varying degrees, but only one algorithm does not lead to severe privacy leakage. Furthermore, existing algorithms fail to meet deployer’s expectations, because they often degrade general model utility and also cannot sustainably accommodate successive unlearning requests or large-scale content removal. Our findings identify key issues with the practicality of existing unlearning algorithms on language models, and we release our benchmark to facilitate further evaluations. [^1] Our dataset and benchmark are available at https://muse-bench.github.io
1 Introduction
Figure 1: MUSE evaluation focuses on six key dimensions of machine unlearning, addressing both data owner and deployer expectations. For example, when an author (data owner) requests the unlearning of the Harry Potter books, they may expect the unlearned model to: (1) avoid generating verbatim copies of the text to protect copyright, (2) eliminate retention of factual knowledge from the books, and (3) not reveal whether the books were previously used in training to protect privacy. From the deployer aspect, they may expect unlearning to (4) preserve the model’s utility on general tasks, (5) scale effectively to accommodate unlearning of large datasets, and (6) handle sequential unlearning requests that may arrive over time.
Training language models (LMs) often involves using vast amounts of text data, which may inadvertently contain private and copyrighted content (Carlini et al., 2021; Henderson et al., 2023; Min et al., 2023; He et al., 2024). In real-world applications, data owners may demand that their data be removed from a trained language model due to privacy or copyright concerns, as mandated for example by the General Data Protection Regulation (GDPR, European Parliament & Council of the European Union, ). Moreover, recent copyright lawsuits (DOE 1 v. GitHub, Inc., N.D. Cal. 2022; Tremblay v. OpenAI, Inc.,, 2023) emphasize the need for removing copyrighted data from the model.
These recent developments have intensified research interest in designing, evaluating, and improving machine unlearning algorithms, which aim to transform an existing trained model into one that behaves as though it had never been trained on certain data (Ginart et al., 2019; Liu et al., 2020; Wu et al., 2020; Bourtoule et al., 2021; Izzo et al., 2021; Gupta et al., 2021; Sekhari et al., 2021; Ye et al., 2022b; Ghazi et al., 2023). Exact unlearning in LMs requires removing the undesired data (the forget set) and retraining the model from scratch on the remaining data (the retain set), which is too costly to be practical, especially for frequent unlearning operations. As such, several efficient approximate unlearning algorithms have been proposed (Eldan & Russinovich, 2023; Zhang et al., 2024b), but existing evaluations of LM unlearning on question answering (Eldan & Russinovich, 2023; Maini et al., 2024) cannot provide a holistic view of how practical and effective a particular unlearning algorithm is. In this work, we propose a systematic, multi-faceted framework called MUSE (Machine Unlearning Six-Way Evaluation; §3) to evaluate six desired properties for unlearning algorithms (Figure 1). Our criteria cover both the data owner’s and the model deployer’s desiderata for a practical unlearning algorithm. Data owners require the LM to unlearn the precise tokens (verbatim memorization), general knowledge encoded in the tokens (knowledge memorization), and any indication that their data was included in the training set to begin with (privacy leakage). On the other hand, model deployers want to effectively accommodate many successive unlearning requests (sustainability) on various sizes of forget sets (scalability) without degrading the general model capabilities (utility preservation).
We apply MUSE to evaluate eight representative machine unlearning algorithms4) on two datasets3.2), focusing on the specific cases of unlearning Harry Potter books and news articles. Our findings indicate that most unlearning algorithms remove verbatim memorization and knowledge memorization with varying degrees of efficacy but operate at the cost of utility preservation and do not effectively prevent privacy leakage (§5.2). In particular, negative preference optimization (NPO; Zhang et al., 2024b) and task vectors (Ilharco et al., 2023) are especially effective in removing these types of memorization, but we find that NPO often permits privacy leakage and both methods induce a sharp drop in the utility of the model. Furthermore, testing their scalability and sustainability reveals that they both algorithms struggle with large forget sets and successive unlearning requests (§5.3).
Our results highlight that unlearning algorithms generally fail to meet data owner expectations in preventing privacy leakage, which is one of the primary motivations for unlearning. Additionally, they struggle to meet all three of the aforementioned deployer expectations. Therefore, although it is increasingly desirable to find an efficient and effective unlearning algorithm amid rising concerns around privacy regulations and copyright litigations, our evaluation suggests that currently feasible unlearning methods are not yet ready for meaningful usage or deployment in real-world scenarios. These findings underscore the pressing need for further research in this area. We also release our benchmark to facilitate further evaluations and welcome extensions to other modalities.
2 Machine Unlearning: Preliminaries and Notations
Machine unlearning (Ginart et al., 2019; Liu et al., 2020; Izzo et al., 2021; Sekhari et al., 2021; Gupta et al., 2021; Ye et al., 2022b; Liu et al., 2024) has emerged as an important capability to accommodate data removal requirements that arise from scenarios with privacy or copyright concerns.
We briefly describe the machine unlearning setting. Consider a dataset $\mathcal{D}_{\textrm{train}}$ and a model $f_{\textrm{target}}$ trained on $\mathcal{D}_{\textrm{train}}$. Suppose we design an algorithm $\mathcal{U}$ to unlearn a specific subset (i.e., the forget set) $\mathcal{D}_{\textrm{forget}}\subset\mathcal{D}_{\textrm{train}}$ from $f_{\textrm{target}}$. We want to preserve performance on a retain set $\mathcal{D}_{\textrm{retain}}=\mathcal{D}_{\textrm{train}}\setminus\mathcal{D}_{\textrm{forget}}$, and we also evaluate the model on an in-distribution but disjoint hold-out set $\mathcal{D}_{\textrm{holdout}}$ which the model has never been trained on. So, the unlearning algorithm $\mathcal{U}$ takes $f_{\textrm{target}}$, $\mathcal{D}_{\textrm{forget}}$, and, optionally, $\mathcal{D}_{\textrm{retain}}$ and outputs an unlearned model $f_{\textrm{unlearn}}$. Exact unlearning ensures $f_{\textrm{unlearn}}$ is behaviorally identical to the model resulting from retraining from scratch, denoted $f_{\textrm{target}}$, but such retraining is usually too costly in real world deployment, so we focus on evaluating approximate unlearning algorithms.
3 The MUSE Evaluation Benchmark
Table 1: Comparison with a previous benchmark: Unlike the previous benchmark TOFU (Maini et al., 2024), which evaluates unlearning on synthetic Q&A datasets, MUSE tackles real-world unlearning challenges: unlearning real-world large-scale corpus (22$\times$ larger) while taking into account six desiderata that are important to both data owners and deployers. More related works are discussed in Appendix 6.
MUSE (ours) TOFU (Maini et al., 2024)
[…]
C2. No knowledge memorization $\checkmark$ $\checkmark$
C3. No privacy leakage $\checkmark$
C4. Utility preservation $\checkmark$ $\checkmark$
C5. Scalability $\checkmark$
C6. Sustainability $\checkmark$
[…]
Data Constitution Verbatim text and knowledge set (Q & A) Q & A
Scale (##\## tokens in forget set) 0.8M for News, 3.3M for Books 0.15M
MUSE evaluates a comprehensive set of desirable properties of machine unlearning across six facets. We detail the evaluation metrics in §3.1 and describe the evaluation corpus in §3.2.
3.1 Evaluation Metrics
Ideally, an unlearned model should behave as if it had never seen the forget set, exhibiting similar behavior to a retrained model on any corpus $\mathcal{D}$ such that $m(f_{\textrm{unlearn}},\mathcal{D})\approx m(f_{\textrm{retrain}},\mathcal{D})$, where $m$ represents any evaluation metric. Prior evaluations on LM unlearning focus on performance of specific tasks like question answering (e.g., Eldan & Russinovich, 2023; Maini et al., 2024). However, these metrics do not faithfully reflect data owner expectations and real-world deployment considerations when performing unlearning. To address this, we propose comprehensive evaluation metrics that consider both data owner and deployer expectations. A comparison between MUSE and the prior benchmark is shown in Table 1.
Data owner expectations. When removing a forget set from a model, data owners typically have three main expectations regarding the unlearned model: (C1) No verbatim memorization: The model should not exactly replicate any details from the forget set. (C2) No knowledge memorization: The model should be incapable of responding to questions about the forget set. (C3) No privacy leakage: It should be impossible to detect that the model was ever trained on the forget set. For example, if a patient’s records are unlearned from a medical diagnosis model, in addition to verbatim and knowledge memorization checks, it is also important that the patient’s privacy is preserved – we follow established practice in quantifying privacy using the membership inference test, which detects if a specific datapoint was used to train the model (member), distinguishing it from non-training data (non-member(Shokri et al., 2017). In this case of unlearning a record from a diagnostic model, it is undesirable for the model to leak membership information, because it would be used to associate the patient with the disease. We quantify these data owner expectations with three evaluation metrics:
C1. No verbatim memorization
When a model has unlearned a medical record, it should not output its contents verbatim. We quantify the verbatim memorization $\mathsf{VerbMem}$ by prompting the model with the first $l$ tokens from a sequence $x_{[:l]}\in\mathcal{D}_{\textrm{forget}}$ and comparing the continuation outputted by the model $f$ to the true continuation $x_{[l+1:]}\in\mathcal{D}_{\textrm{forget}}$ using the ROUGE-L F1 score (Lin, 2004).
$\mathsf{VerbMem}(f,\mathcal{D}):=\frac{1}{|\mathcal{\mathcal{D}_{\textrm{forget}}}|}\sum_{x\in\mathcal{\mathcal{D}_{\textrm{forget}}}}\mathsf{ROUGE}(f(x_{[:l]}),x_{[l+1:]})$
C2. No knowledge memorization
When a model has unlearned a medical record, it should no longer be able to answer questions about that record. We measure a model $f$’s memorization of knowledge from the forget set $\mathcal{\mathcal{D}_{\textrm{forget}}}$ as follows: for each example $x\in\mathcal{\mathcal{D}_{\textrm{forget}}}$ associated with a question-answer pair $(q,a)$, [^2] Examples of question-answer pairs derived from the original corpus can be found in Appendix D. we gather the model’s answer to the question $q$, denoted $f(q)$. We then average the ROUGE scores for all question-answer pairs in $\mathcal{\mathcal{D}_{\textrm{forget}}}$ to compute the knowledge memorization score $\mathsf{KnowMem}$:
$\mathsf{KnowMem}(f,\mathcal{\mathcal{D}_{\textrm{forget}}}):=\frac{1}{|\mathcal{\mathcal{D}_{\textrm{forget}}}|}\sum_{(q,a)\in\mathcal{\mathcal{D}_{\textrm{forget}}}}\mathsf{ROUGE}(f(q),a)$
Figure 2: Distribution of the MIA metric (see C3) for $\mathcal{D}_{\textrm{forget}}$, $\mathcal{D}_{\textrm{holdout}}$, and $\mathcal{D}_{\textrm{retain}}$. Differences in the metric between forget and holdout sets indicate various unlearning outcomes of $\mathcal{D}_{\textrm{forget}}$, potentially leaking privacy. A perfectly unlearned model (b) should show similar MIA metrics distribution for $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{holdout}}$. Unlearning methods may fail by under-unlearning $\mathcal{D}_{\textrm{forget}}$, making it similar to $\mathcal{D}_{\textrm{retain}}$ (c), or over-unlearning it, causing divergence from $\mathcal{D}_{\textrm{holdout}}$ (d).
C3. No privacy leakage
As discussed previously, it is desirable that the unlearned model does not leak membership information indicating that $\mathcal{D}_{\textrm{forget}}$ was part of $\mathcal{D}_{\textrm{train}}$. To determine if a given example was used during training, membership inference attack (MIA) exploits distributional differences in certain statistics (e.g., loss) between training (member) and non-training (non-member) data: if the loss on the example is low, then it was likely used for training. As shown in Figure 2, unlearning typically increases the loss on the example, but there are two possible ways that unlearning can fail to prevent privacy leakage: (1) under-unlearning, when the loss is not made large enough; and (2) over-unlearning, when the loss is made abnormally large. To accurately measure the privacy leakage, we employ Min-KProb (Shi et al., 2024a) , a state-of-the-art MIA method for LMs based on the loss, and compute the standard AUC-ROC score (Murakonda et al., 2021; Ye et al., 2022a) of discriminating $\mathcal{D}_{\textrm{forget}}$ (members) and $\mathcal{D}_{\textrm{holdout}}$ (non-members). [^3] An MIA algorithm compares its score to a given threshold to classify a given datapoint as a member or non-member. The AUC-ROC is a single value that summarizes the overall performance of the MIA algorithm by measuring its ability to discriminate between members and non-members across all possible thresholds. By comparing the AUC score with that of the retrained model, we define [^4] Generally, $\textsf{AUC}(f_{\textrm{retrain}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})\approx 0.5$, though sometimes there are intrinsic distribution shifts between $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{holdout}}$ that may bias the baseline away from 0.5.
$\textsf{PrivLeak}:=\frac{\textsf{AUC}(f_{\textrm{unlearn}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})-\textsf{AUC}(f_{\textrm{retrain}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})}{\textsf{AUC}(f_{\textrm{retrain}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})},$
The PrivLeak metric for a good unlearning algorithm should be close to zero, whereas an over/under-unlearning algorithm will get a large positive/negative metric.
Deployer expectations. Model deployers have their own considerations for using unlearning algorithms in the real world. Unlearning specific datapoints can unpredictably degrade model capabilities in ways that are difficult to recover. Moreover, deployers are expected to effectively accommodate somewhat large-scale forget sets and successive unlearning requests from data owners. As such, we consider three key metrics: (C4) utility preservation on the retain set, (C5) scalability to handle large-scale content removal, and (C6) sustainability to maintain performance over sequential unlearning requests.
C4. Utility preservation.
Model capabilities are often hard-won through expensive training procedures, so deployers would want an unlearning algorithm that preserves performance on the retain set. To quantify this, we evaluate the unlearned model’s performance on the retain set using the knowledge memorization metric $\mathsf{KnowMem}(f_{\textrm{unlearn}},\mathcal{D}_{\textrm{retain}})$.
C5. Scalability.
We assess the scalability of unlearning methods by examining their performance on forget sets of varying sizes. Let $\mathcal{D}_{u}^{c}$ denote a forget set of size $c$, and $f_{u}^{c}$ be the corresponding unlearned model. For any data owner-valued metric such as utility preservation, we measure scalability by analyzing the trend of this metric as $c$ increases from small to large values.
C6. Sustainability.
Machine unlearning operations often need to be applied sequentially, as data removal requests may arrive at different times. [^5] For example, under GDPR, if Alice requests the removal of her data and Bob submits another removal request 31 days later, both requests must be fulfilled within 30 days. This requires the model deployer to first unlearn Alice’s data and then process Bob’s request on the updated model. We denote the unlearned model after processing the $k$-th request as $f_{u,k}$. To measure sustainability, we analyze the trend of any data owner-valued metric as the number of sequential unlearning requests $k$ increases.
3.2 Evaluation Corpus
MUSE considers two representative types of textual data that may frequently involve unlearning requests: news articles (Tremblay v. OpenAI, Inc.,, 2023) and books (Eldan & Russinovich, 2023). These datasets are detailed as follows:
  • News consists of BBC news articles (Li et al., 2023b) collected after August 2023. All articles are randomly divided into (disjoint) forget, retain, and holdout sets.
  • Books consists of the Harry Potter book series. To simulate a real-world setting for testing utility preservation (C4), we include different types of materials in the forget and retain sets. The forget set contains the original books, while the retain set contains related content from the Harry Potter FanWiki,666harrypotter.fandom.com/wiki representing domain knowledge that should be retained after unlearning.
Table 2: Examples of MUSE. Each corpus has Verbatim text and Knowledge sets (QA pairs derived from the original text) for evaluating verbatim and knowledge memorization. In News, $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{retain}}$ are two disjoint sets of news articles. In Books, $\mathcal{D}_{\textrm{forget}}$ is the Harry Potter book series while $\mathcal{D}_{\textrm{retain}}$ consists of wiki articles about the series. The sizes of the forget and retain sets are reported in tokens in ().
Corpus Forget Set Retain Set
News Article (0.8 M tokens) News Article (1.6 M tokens)
[…]
A: A father whose 12-year-old son
Harry Potter Books (1.1 M tokens) Harry Potter FanWiki (0.5 M tokens)
[…]
Portuguese for ‘open’.
[…]
A: Portuguese
For each corpus, we construct: 1) Verbatim text: the original text to assess the unlearning methods to remove verbatim memorization (C1), and 2) Knowledge set: a set of derived (question, answer) pairs based on the original texts to evaluate the unlearning method’s effectiveness in purging learned knowledge and preventing knowledge memorization (C2). To create the Knowledge set, we partition the Verbatim text into excerpts and use GPT-4 (OpenAI, 2023) to generate (question, answer) pairs for each excerpt. For more details about the dataset generation pipeline, see Appendix D.
Appendix D provides examples from the news and books corpora. The details of the dataset splits and dataset sizes are provided in Appendix D.
4 Unlearning Methods
We evaluate eight efficient approximate unlearning methods belonging to four families of algorithms.
Four families of unlearning methods. We first introduce four families of unlearning methods, which serve as the basis for the eight methods we evaluate.
  • Gradient Ascent (GA) minimizes the likelihood of correct predictions on $\mathcal{D}_{\textrm{forget}}$ by performing gradient ascent on the cross-entropy loss (the opposite of conventional learning with gradient descent). GA has achieved mixed results: while Jang et al. (2023) found it effective for unlearning examples from the Enron email dataset (Klimt & Yang, 2004) with minimal performance degradation, Ilharco et al. (2023) reported that GA significantly harms general model utility when unlearning a high-toxicity subset of the Civil Comments dataset (Borkan et al., 2019).
  • Negative Preference Optimization (NPO; Zhang et al., 2024b) treats the forget set as negative preference data and adapts the offline DPO objective (Rafailov et al., 2023) to tune the model to assign low likelihood to the forget set without straying too far from the original model $f_{\textrm{target}}$.
    $\mathcal{L}_{\mathrm{NPO}}(\theta)=-\frac{2}{\beta}\mathbb{E}_{x\sim\mathcal{D}_{\textrm{forget}}}\left[\log\sigma\left(-\beta\log\frac{f_{\theta}(x)}{f_{\textrm{target}}(x)}\right)\right],$
    where $f_{\theta}$ refers to the model that undergoes unlearning, $\sigma$ is the sigmoid function, and $\beta$ is a hyperparameter that controls the allowed divergence of $f_{\theta}$ from its initialization $f_{\textrm{target}}$. Following Rafailov et al. (2023); Zhang et al. (2024b), we fix $\beta=0.1$ in our experiments.
  • Task Vectors (Ilharco et al., 2023) derived from straightforward arithmetic on the model weights can effectively steer neural network behavior. We adapt task vectors to perform unlearning in two stages. First, we train $f_{\textrm{target}}$ on $\mathcal{D}_{\textrm{forget}}$ until the model overfits, yielding a reinforced model $f_{\textrm{reinforce}}$. We then obtain a task vector related to $\mathcal{D}_{\textrm{forget}}$ by calculating the weight difference between $f_{\textrm{target}}$ and $f_{\textrm{reinforce}}$. To achieve unlearning, we subtract this task vector from $f_{\textrm{target}}$’s weights, intuitively moving the model away from the direction it used to adapt to $\mathcal{D}_{\textrm{forget}}$ – i.e., $f_{\textrm{unlearn}}=f_{\textrm{target}}-(f_{\textrm{reinforce}}-f_{\textrm{target}})$.
  • Who’s Harry Potter (WHP; Eldan & Russinovich, 2023) defines the unlearned model $f_{\textrm{unlearn}}$ as the interpolation between the target model $f_{\textrm{target}}$ and the reinforced model $f_{\textrm{reinforce}}$. Let $p_{f}(\cdot|x)$ denote the token distribution parametrized by the model $f$ when given a prompt $x$ as input. Then, concretely, for any input $x$, WHP samples the next token from p_f_unlearn(⋅| x) = p_f_target(⋅| x) - α(p_f_reinforce(⋅| x) - p_f_target(⋅| x)) where $\alpha$ is a hyperparameter that controls the interpolation between the two models.
Two regularizers for utility preservation. GA and NPO are not explicitly designed for utility preservation, so we discuss several regularization strategies that either improve the performance on the retain set or ensure the unlearned model remains close to the target model during unlearning.
  • Gradient Descent on the Retain Set (GDR; Liu et al., 2022; Maini et al., 2024; Zhang et al., 2024b) augments the unlearning objective with a standard gradient descent learning objective on the cross-entropy of the retain set $\mathcal{D}_{\textrm{retain}}$ to more directly train the model to maintain its performance on $\mathcal{D}_{\textrm{retain}}$.
  • KL Divergence Minimization on the Retain Set (KLR; Maini et al., 2024; Zhang et al., 2024b) encourages the unlearned model’s probability distribution $p_{f_{\textrm{unlearn}}}(\cdot|x)$ to be close to the target model’s distribution $p_{f_{\textrm{target}}}(\cdot|x)$ on inputs from the retain set $x\in\mathcal{D}_{\textrm{retain}}$.
List of methods. We combine GA and NPO with the two regularizers GDR and KLR,777 These regularizers are not compatible with Task Vector and WHP, because Task Vector involves purposefully overfitting a model to $\mathcal{D}_{\textrm{forget}}$ when deriving the task vector, and WHP is a test-time technique where the unlearning operation involves no optimization by itself. which yields four new combinations. Hence, we end up with a total of 8 candidate unlearning methods: GA, $\textsf{GA}_{\textsf{GDR}}$, $\textsf{GA}_{\textsf{KLR}}$, NPO, $\textsf{NPO}_{\textsf{GDR}}$, $\textsf{NPO}_{\textsf{KLR}}$, Task Vector, and WHP. In general, the cost of the approximate unlearning method is negligible compared to retraining. Details about the efficiency of these methods are reported in Appendix B.3.
5 Experiments
We evaluate the eight representative unlearning methods using the experimental setup described in §5.1. We present the results for data owner expectations in §5.2 and for deployer expectations in §5.3.
5.1 Experimental Setup
Table 3: Most unlearning methods effectively remove verbatim and knowledge memorization but significantly impact utility and privacy. We evaluate the 8 algorithms described in §4 on 4 of the criteria in MUSE. We include the results of $f_{\textrm{retrain}}$ for reference. We highlight results in blue if the unlearning algorithm satisfies the criterion and highlight it in orange otherwise. For privacy leakage, large positive values suggest   over-unlearning, while large negative values suggest   under-unlearning (see §3.1). This table covers the results for C1 to C4, while results for C5 and C6 are shown in Figure 6.
C1. No Verbatim Mem. C2. No Knowledge Mem. C3. No Privacy Leak. C4. Utiltiy Preserv.
VerbMem on $\mathcal{D}_{\textrm{forget}}$ ($\downarrow$) KnowMem on $\mathcal{D}_{\textrm{forget}}$ ($\downarrow$) PrivLeak ($\in[-5\%,5\%]$) KnowMem on $\mathcal{D}_{\textrm{retain}}$ ($\uparrow$)
News
Target $f_{\textrm{target}}$ $58.4$ $63.9$ $-99.8$ $55.2$
Retrain $f_{\textrm{retrain}}$ $\mathbf{20.8}$ $\mathbf{33.1}$ $\mathbf{0.0}$ $\mathbf{55.0}$
GA $0.0$downarrow1000.00.$0$downarrow1005.25.25.25.2 over-unlearn $0.0$downarrow100%
$\textsf{GA}_{\textsf{GDR}}$ $4.9$downarrow76.531.031.031.031.0 $\downarrow$6.3108.1108.1108.1108.1 over-unlearn $27.3$downarrow50.3%
$\textsf{GA}_{\textsf{KLR}}$ $27.4$uparrow31.450.250.250.250.2 $\uparrow$51.596.196.1-96.1- 96.1 under-unlearn $44.8$downarrow18.5%
NPO $0.0$downarrow1000.00.$0$downarrow10024.424.424.424.4 over-unlearn $0.0$downarrow100.0%
$\textsf{NPO}_{\textsf{GDR}}$ $1.2$downarrow94.454.654.654.654.6 $\uparrow$64.8105.8105.8105.8105.8 over-unlearn $40.5$downarrow26.3%
$\textsf{NPO}_{\textsf{KLR}}$ $26.9$uparrow29.049.049.049.049.0 $\uparrow$48.195.895.8-95.8- 95.8 under-unlearn $45.4$downarrow17.4%
Task Vector $57.2$uparrow174.766.266.266.266.2 $\uparrow$100.099.899.8-99.8- 99.8 under-unlearn $55.8$uparrow1.5%
WHP $19.7$downarrow5.621.221.221.221.2 $\downarrow$35.9109.6109.6109.6109.6 under-unlearn $28.3$downarrow48.5%
Books
Target $f_{\textrm{target}}$ $99.8$ $59.4$ $-57.5$ $66.9$
Retrain $f_{\textrm{retrain}}$ $\mathbf{14.3}$ $\mathbf{28.9}$ $\mathbf{0.0}$ $\mathbf{74.5}$
GA $0.0$downarrow1000.00.$0$downarrow10025.025.0-25.0- 25.0 under-unlearn $0.0$downarrow100%
$\textsf{GA}_{\textsf{GDR}}$ $0.0$downarrow1000.00.$0$downarrow10026.526.5-26.5- 26.5 under-unlearn $10.7$downarrow85.6%
$\textsf{GA}_{\textsf{KLR}}$ $16.0$uparrow11.421.921.921.921.9 $\downarrow$24.440.240.2-40.2- 40.2 under-unlearn $37.2$downarrow50.0%
NPO $0.0$downarrow1000.00.$0$downarrow10024.324.3-24.3- 24.3 under-unlearn $0.0$downarrow100%
$\textsf{NPO}_{\textsf{GDR}}$ $0.0$downarrow1000.00.$0$downarrow10030.830.8-30.8- 30.8 under-unlearn $22.8$downarrow69.4%
$\textsf{NPO}_{\textsf{KLR}}$ $17.0$uparrow18.225.025.025.025.0 $\downarrow$13.443.543.5-43.5- 43.5 under-unlearn $44.6$downarrow40.1%
Task Vector $99.7$uparrow595.052.452.452.452.4 $\uparrow$81.257.557.5-57.5- 57.5 under-unlearn $64.7$downarrow13.1%
WHP $18.0$uparrow25.255.755.755.755.7 $\uparrow$92.956.556.556.556.5 over-unlearn $63.6$downarrow14.6%
[…]
Unlearning experimental configuration. Following prior work (Maini et al., 2024), we run GA, NPO, and their regularized variants using the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of $10^{-5}$ and a batch size of 32. We employ the stopping criteria as follows: if the utility (i.e., KnowMem on $\mathcal{D}_{\textrm{retain}}$) of a model undergoing unlearning drops below that of $f_{\textrm{retrain}}$ within 10 epochs of unlearning, we stop at the first epoch where this condition holds; otherwise, we take a checkpoint from the 10th epoch. For Task Vector and WHP, to obtain the reinforced model for unlearning, we fine-tune the target model for 10 epochs using the same learning rate and batch size. Further details on the model fine-tuning and unlearning can be found in Section B.2.
5.2 Results: Data Owner Expectations
We first analyze how eight unlearning methods meet data owner expectations (C1, C2 & C3 in §3.1).
C1&C2. Most methods are effective for unlearning memorization. As shown in Table 3, most unlearning methods perform exceptionally well in [C1. No verbatim memorization] and [C2. No knowledge memorization], often reducing VerbMem and KnowMem even beyond the levels achieved by the retrained model. Notably, some methods, such as GA and NPO, achieve a score of 0 for both VerbMem and KnowMem, meaning that these methods completely prevent the unlearned models from producing any text related to the forget set. However, as we will see later, these reductions often come at the cost of significant utility loss on the retain set.
Figure 3: Distribution of Min-KProb, an MIA metric, for $\mathcal{D}_{\textrm{forget}}$, $\mathcal{D}_{\textrm{holdout}}$, and $\mathcal{D}_{\textrm{retain}}$. Consistent with the expected pattern in Figure 2, $f_{\textrm{retrain}}$ shows perfect unlearning, with the overlapping distributions for $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{holdout}}$. Existing approximate unlearning methods typically either under-unlearn or over-unlearn. For example, $\textsf{GA}_{\textsf{KLR}}$ shows slight under-unlearning, while $\textsf{GA}_{\textsf{GDR}}$ over-unlearns, pushing the Min-KProb of $\mathcal{D}_{\textrm{forget}}$ to an extreme level.
[…]
Figure 5: Utility preservation vs. knowledge memorization on BBC. $f_{\textrm{retrain}}$ maintains high utility on $\mathcal{D}_{\textrm{retain}}$ while showing low knowledge memorization on $\mathcal{D}_{\textrm{forget}}$. GA and NPO without regularizers show significant utility loss, collapsing to the origin. Every other unlearning method unlearns the knowledge on $\mathcal{D}_{\textrm{forget}}$ at the cost of utility.
C3. Unlearning leads to privacy leakage. Most unlearning methods reveal the membership of $\mathcal{D}_{\textrm{forget}}$ in $\mathcal{D}_{\textrm{train}}$ through under-unlearning ($\textsf{PrivLeak}\ll 0$) or over-unlearning ($\textsf{PrivLeak}\gg 0$), as shown in Table 3. We further examine the effectiveness of membership inference by plotting ROC curves in Figure 5. The deviation from the diagonal line indicates the attacker’s advantage over random guessing. We observe that the Min-KProb based attack achieves $\textsf{AUC}\approx 0$ on $f_{\textrm{target}}$, confirming its effectiveness. Meanwhile, the ROC curve for $f_{\textrm{retrain}}$ closely follows the diagonal line ($\textsf{AUC}=0.47$), suggesting that perfect unlearning ensures MIA is no more effective than random guessing. Among the approximate unlearning methods, GA and $\textsf{NPO}_{\textsf{GDR}}$ without regularizers consistently over-unlearn ($\textsf{AUC}>0.7$), whereas KLR-regularized methods ($\textsf{NPO}_{\textsf{KLR}}$ and $\textsf{GA}_{\textsf{KLR}}$) tend to under-unlearn and barely improve privacy leakage over $f_{\textrm{target}}$. WHP also deviates from the diagonal significantly.
In Figure 3, we further visualize the distribution of Min-KProb, the MIA metric computed across $\mathcal{D}_{\textrm{forget}}$, $\mathcal{D}_{\textrm{retain}}$, and $\mathcal{D}_{\textrm{holdout}}$. The behavior of $f_{\textrm{target}}$ and $f_{\textrm{retrain}}$ mirrors the patterns sketched in Figure 2, where $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{retain}}$ are distinguishable in $f_{\textrm{target}}$ but overlap in $f_{\textrm{retrain}}$. Existing approximate unlearning methods typically either under-unlearn or over-unlearn. For example, $\textsf{GA}_{\textsf{KLR}}$ does not sufficiently increase the Min-KProb metric for $\mathcal{D}_{\textrm{forget}}$ to align with the distribution of $\mathcal{D}_{\textrm{holdout}}$, indicating under-unlearning. On the other hand, $\textsf{NPO}_{\textsf{GDR}}$ over-unlearns, significantly raising the MIA metric across all datasets and especially for $\mathcal{D}_{\textrm{forget}}$.
5.3 Results: Deployment Considerations
C4. Unlearning significantly degrades model utility. Table 3 [C4 Utility Preserv.] shows that all unlearning methods compromise the model’s utility by $24.2\%\sim 100\%$. Notably, several methods (GA, $\textsf{GA}_{\textsf{GDR}}$, $\textsf{NPO}_{\textsf{GDR}}$) lead to complete utility loss, rendering the unlearned models practically unusable. Figure 5 illustrates the trade-offs between utility preservation on $\mathcal{D}_{\textrm{retain}}$ and knowledge memorization on $\mathcal{D}_{\textrm{forget}}$. An ideal unlearned model should mimic the behavior of $f_{\textrm{retrain}}$ (desired region) by achieving a low level of memorization on $\mathcal{D}_{\textrm{forget}}$ while maintaining its utility. However, most methods, such as $\textsf{GA}_{\textsf{KLR}}$, $\textsf{NPO}_{\textsf{KLR}}$, and WHP, unlearn the knowledge on $\mathcal{D}_{U}$ at the cost of utility.
Figure 6: The performance of GA, NPO, and their regularized variants, measured by utility preservation, degrades with larger forget set sizes (a) and sequential unlearning requests (b).
C5. Unlearning methods scale poorly with forget set sizes. To evaluate the robustness of the unlearning methods to larger forget sets, we collect additional news articles from the same distribution to scale our News corpus from 0.8M tokens to 3.3M tokens and observe the utility preservation at four different forget set sizes. As shown in Figure 6 (a), the model utility decrease with the size of the forget set and achieves a minimum at the largest size.
C6. Unlearning methods cannot sustainably accommodate sequential unlearning requests. To evaluate the robustness of these unlearning methods to more than one unlearning requests, we sequentially apply $k$ unlearning processes, each with respect to a different forget set. To simulate sequential unlearning, we partition the extended News forget set (comprised of 3.3M tokens) into four disjoint folds (each containing 0.8M tokens) and apply the unlearning methods to each fold in a sequential manner.
We again select utility preservation as the target metric for comparison. As shown in Figure 6 (b), the performance of an unlearned model tends to decrease significantly with respect to the number of unlearning requests, indicating that current unlearning methods are not yet ready to handle sequential unlearning in a sustainable manner.
6 Related Work
Machine unlearning for non-language model applications.
Machine unlearning is a long-running, well-studied topic. Several studies have explored exact unlearning, aiming to make the unlearned model ($f_{\textrm{unlearn}}$) exactly identical to the reference model ($f_{\textrm{retrain}}$). As expected, this can only be accomplished in simple models like SVMs (Cauwenberghs & Poggio, 2000; Tveit et al., 2003; Romero et al., 2007; Karasuyama & Takeuchi, 2010) or naive Bayes models (Cao & Yang, 2015). Another approach is to ensure that the unlearned model $f_{\textrm{unlearn}}$ is probabilistically indistinguishable from $f_{\textrm{retrain}}$ (Ginart et al., 2019; Guo et al., 2020), and this view of certifiable unlearning is closely related to differential privacy (Dwork et al., 2006b, a). This rigorous definition of unlearning has inspired several theoretical works that characterize the feasibility of unlearning in convex and non-convex models, but those proposed algorithms are too computationally costly to operate on modern-day LMs (Izzo et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Gupta et al., 2021). Several more tractable unlearning algorithms have been proposed (Borkan et al., 2019; Ginart et al., 2019; Thudi et al., 2022; Chourasia & Shah, 2023) with broader applications such as image classification (Ginart et al., 2019; Golatkar et al., 2020a), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023; Fan et al., 2023), Federated Learning (Liu et al., 2020; Che et al., 2023; Halimi et al., 2022; Huang et al., 2022) and Recommender Systems (Li et al., 2024b).
Machine unlearning for language models: methods and applications.
Machine unlearning has recently found its way into language model applications. In §4, we discuss some standard unlearning methods based on parameter optimization, like the Gradient Ascent and its variance. Other notable non-training-based unlearning methods include localization-informed unlearning (Meng et al., 2022; Wu et al., 2023; Wei et al., 2024a), which involves identifying model units (e.g., layers, neurons) closely related to the unlearning data or tasks and then locally editing and modifying the units. In-context unlearning (Pawelczyk et al., 2023) offers another approach, treating the model as a black box and modifying its output results using external knowledge.
Machine unlearning has also been applied to various downstream language model tasks, though the unit of machine unlearning may differ from what we study in this work. Our evaluation focuses on unlearning specific examples or datasets, aiming to make LMs forget either the phrasing or the content knowledge of targeted data, while preserving their utility for data not targeted for removal. This is crucial for ensuring privacy and copyright compliance. In addition to this specific unlearning, there’s also a broader application similar to model editing, where outdated information is replaced with new knowledge (Pawelczyk et al., 2023; Yu et al., 2023; Belrose et al., 2024). Moreover, efforts have been made to eliminate harmful behaviors in language models by creating toxicity benchmarks and enhancing safety measures (Lu et al., 2022; Yao et al., 2023; Li et al., 2024a; Zhang et al., 2024b). Despite these varied approaches to unlearning at different operational and knowledge levels, the evaluation principles we propose such as preserving utility, ensuring scalability, and maintaining sustainability—are relevant across these contexts.
Machine unlearning for language models: evaluation.
Evaluating machine unlearning methods for language model applications is also critical. Most previous studies have focused this evaluation on specific tasks such as question answering or sentence completion. For example, Eldan & Russinovich (2023) experiment with unlearning to forget Harry Potter books and demonstrate the effectiveness of their methods by showing that familiarity scores, measured through completion-based, token-probability-based, and question-answering evaluations, significantly decline post-unlearning. Lynch et al. (2024) further suggest comparing unlearned models with perfectly retrained models. Their evaluation finds that while familiarity scores with the forget set may drop post-unlearning, they still remain higher than those of the retrained model. Wei et al. (2024b) evaluate the feasibility of using unlearning techniques to prevent language models from generating copyrighted content. The closest work to ours is TOFU (Maini et al., 2024), a benchmark featuring 200 synthetic author profiles, each with 20 question-answer pairs, divided into forget and retain sets. However, TOFU is relatively small-scale (0.15M tokens) and focuses on the evaluation of question answering. Additionally, current evaluations focus on limited aspects of data owner expectations and do not adequately reflect real-world deployment considerations, such as scalability and potential sequential unlearning requests. In contrast, MUSE formally defines different unlearning scopes and corresponding metrics, resulting in a systematic six-way evaluation featuring both data owners’ and deployers’ expectations. The evaluation uses a large-scale corpus of over 6 million tokens, separated into verbatim text and knowledge sets. We also note that some of our findings align with previous evaluations. For example, our observation that over- or under-unlearn can exacerbate privacy leakage (§5.2) is consistent with the recent work by Hayes et al. (2024). Our findings align with the the concurrent study by Shumailov et al. (2024) showing that unlearning gives a false sense of security as unlearned knowledge can resurface through in-context learning.
Survey papers.
We direct readers to several insightful survey papers for further reading. For non-LLM applications, notable surveys include Shintre et al. (2019); Nguyen et al. (2022); Thudi et al. (2022); Xu et al. (2023). Additionally, the NeurIPS 2023 machine unlearning competition for image classification [^8] https://unlearning-challenge.github.io is a valuable source of empirical methods tailored for this specific application (Triantafillou et al., 2023). For language model applications, Si et al. (2023) categorize unlearning methods into different families and summarize datasets for evaluating unlearning. Liu et al. (2024) review LM unlearning algorithms by targets and methods, discuss the effectiveness and efficiency of existing approaches and emphasize the importance of clearly defining the unlearning scope.
7 Conclusion
In this work, we propose MUSE, a comprehensive machine unlearning evaluation benchmark that highlights six desirable properties from the perspectives of both data owners and model deployers. We find that current unlearning methods successfully prevent the model’s memorization of content at a significant cost to utility on data not intended for removal. They also lead to severe privacy leakage and cannot sustainably accommodate successive unlearning requests or large-scale content removal. These findings highlight the need for future research into more robust unlearning methods.
Limitations. While MUSE provides a systematic benchmark for evaluating unlearning algorithms, it does not consider all possible considerations. For example, data owners may have additional expectations, such as ensuring their information cannot be probed from intermediate activations (Song & Raghunathan, 2020) or receiving formal guarantees of unlearning success (Sekhari et al., 2021; Gupta et al., 2021; Ghazi et al., 2023). Similarly, deployers may expect other capabilities, like fine-tuning and in-context learning, to be preserved, and may prefer unlearning algorithms that are both computationally efficient and storage-wise cheap (e.g. does not need to keep a copy of the retain set). MUSE currently evaluates unlearning for language models using books and news articles, but it could be extended to other corpora, such as medical notes (Johnson et al., 2016, 2020) and emails (Klimt & Yang, 2004), which often involve privacy concerns (Li et al., 2023a; Huang et al., 2023). We also plan to evaluate different-sized LMs in the future. Finally, our approach can be generalized to construct multi-faceted benchmarks for multimodal models (Golatkar et al., 2020b; Cheng & Amiri, 2023; Zhang et al., 2024c). Further discussion on broader impact are in Appendix A.
8 Acknowledgements
We thank Eric Wallace, Robin Jia, Howard Chen, and anonymous reviewers of the GenLaw workshop for the valuable feedback and discussions.
Appendix A Broader Impact
As LMs are deployed broadly and publicly, there is mounting legal and social pressure on deployers to release models that permit effective unlearning when requested by data owners (European Parliament & Council of the European Union, ; DOE 1 v. GitHub, Inc., N.D. Cal. 2022; Tremblay v. OpenAI, Inc.,, 2023). These incentives have prompted a flurry of new unlearning algorithms stemming from different technical perspectives. As such, systematic evaluation of the strengths and weaknesses of these methods when executing realistic unlearning requests on popular models is essential. MUSE disentangles several desirable properties of unlearning algorithms and finds that no existing algorithm is able to satisfy all of the data owner and deployer considerations. We hope that our fine-grained, multi-faceted framework facilitates the improvement of unlearning algorithms. Moreover, we expect that the general approach of designing metrics to balance the considerations of various stakeholders is flexible and can adapt to the rapidly shifting legal, social, and economic landscape.
We also acknowledge the potential negative impacts of our study. One limitation of our evaluation benchmark is that we do not have comprehensive study of how unlearning would impact the model performance for different user bases, especially underrepresented groups. However, we note proper handling and evaluation of fairness issues in unlearning is still an active ongoing research area (Zhang et al., 2024a; Oesterling et al., 2024), therefore we leave it as future work. Additionally, our work may be misinterpreted towards skepticism regarding the broader use of machine unlearning, as our current evaluation reveals that existing unlearning methods are not yet ready for effective real-world deployment. However, machine unlearning, especially for large language models, is a young and active research area and new algorithms are constantly being proposed. We emphasize that our results is not a criticism of the paradigm of machine unlearning, but a study of the potential downsides of existing methods and a call for better algorithms. We believe our benchmark is an important step towards guiding future algorithm design of machine unlearning research towards more realistic deployment scenarios.
Appendix B Experimental Details
B.1 Compute Configurations
All experiments are conducted on 8 NVIDIA A40 GPU cards in a single node.
B.2 Experimental Setup
Finetuning details.
[…]
Unlearning details.
For all the unlearning methods in Table 3, we use a constant learning rate of $10^{-5}$ and a batch size of 32. For $f_{\text{reinforced}}$ used in WHP and Task Vector, we fine-tune $f_{\textrm{target}}$ for 10 epochs.
Before evaluation, for each unlearning method, we select its optimal epoch or $\alpha$ (both of which are parameters that control a degree of unlearning) by using our unlearning stopping criteria based on the unlearned model’s utility on $\mathcal{D}_{\textrm{retain}}$ compared to that of $f_{\textrm{retrain}}$. The chosen epochs or $\alpha$’s for each method are listed below.
Table 4: Optimal epochs or $\alpha$’s for each unlearning method.
Unlearning Method News Books
GA epoch 1 epoch 1
$\textsf{GA}_{\textsf{GDR}}$ epoch 7 epoch 1
$\textsf{GA}_{\textsf{KLR}}$ epoch 10 epoch 5
NPO epoch 1 epoch 1
$\textsf{NPO}_{\textsf{GDR}}$ epoch10 epoch 1
$\textsf{NPO}_{\textsf{KLR}}$ epoch 10 epoch 4
Task Vector $\alpha=2^{9}$ $\alpha=2^{9}$
WHP $\alpha=2^{2}$ $\alpha=2^{8}$
B.3 Efficiency of Unlearning Methods
We report the efficiency of unlearning methods in Table 5, measured by the wall-clock time for a single gradient update step of unlearning. The time measurements were conducted using 8 NVIDIA A40 GPUs on a single node, with a batch size of 32 and an input length of 2048 tokens. Each step corresponds to one gradient update processing a total of 65,536 tokens (32 $\times$ 2048 tokens). For Task Vector and WHP, each step represents one iteration of fine-tuning to create the reinforced model.
Table 5: Wall-clock time required for each unlearning method, measured in seconds per step.
Unlearning Method Time (Seconds/Step)
GA 4.14
$\textsf{GA}_{\textsf{GDR}}$ 6.05
$\textsf{GA}_{\textsf{KLR}}$ 7.58
NPO 5.68
$\textsf{NPO}_{\textsf{GDR}}$ 7.59
$\textsf{NPO}_{\textsf{KLR}}$ 9.11
Task Vector 4.14
WHP 4.14
Appendix C More Experimental Results
C.1 Confidence Intervals for C1, C2 and C4 in Table 3
We compute confidence intervals for C1, C2, and C4 (Mean ROUGE-L F1) using bootstrapping [^9] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.bootstrap.html. For each mean ROUGE-L score reported in Table 3, we draw 9,999 bootstrap resamples and calculate a two-tailed 95% confidence interval using the “percentage” method.
Table 6: 95% confidence intervals computed for mean Rouge-L scores used in C1, C2, and C4.
C1. No Verbatim Mem. C2. No Knowledge Mem. C4. Utiltiy Preserv.
VerbMem on $\mathcal{D}_{\textrm{forget}}$ ($\downarrow$) KnowMem on $\mathcal{D}_{\textrm{forget}}$ ($\downarrow$) KnowMem on $\mathcal{D}_{\textrm{retain}}$ ($\uparrow$)
News
Target $f_{\textrm{target}}$ $58.4$ [54.1, 62.9] $63.9$ [58.7, 69.0] $55.2$ [50.7, 59.9]
Retrain $f_{\textrm{retrain}}$ $\mathbf{20.8}$ [18.5, 23.7] $\mathbf{33.1}$ [26.8, 39.5] $\mathbf{55.0}$ [50.3, 59.8]
GA $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0]
$\textsf{GA}_{\textsf{GDR}}$ $4.9$ [4.5, 5.2] $31.0$ [24.2, 38.0] $27.3$ [21.9, 33.0]
$\textsf{GA}_{\textsf{KLR}}$ $27.4$ [25.1, 29.9] $50.2$ [43.1, 56.9] $44.8$ [39.2, 50.5]
NPO $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0]
$\textsf{NPO}_{\textsf{GDR}}$ $1.2$ [0.3, 2.3] $54.6$ [47.5, 61.5] $40.5$ [34.7, 46.2]
$\textsf{NPO}_{\textsf{KLR}}$ $26.9$ [24.7, 29.3] $49.0$ [41.8, 61.5] $45.4$ [39.8, 51.1]
Task Vector $57.2$ [52.6, 62.0] $66.2$ [61.3, 71.2] $55.8$ [51.0, 60.6]
WHP $19.7$ [17.8, 21.6] $21.2$ [16.0, 26.7] $28.3$ [23.3, 33.4]
Books
Target $f_{\textrm{target}}$ $99.8$ [99.8, 99.9] $59.4$ [52.7, 66.0] $66.9$ [59.6, 73.8]
Retrain $f_{\textrm{retrain}}$ $\mathbf{14.3}$ [13.6, 15.1] $\mathbf{28.9}$ [22.1, 35.7] $\mathbf{74.5}$ [68.4, 80.0]
GA $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0]
$\textsf{GA}_{\textsf{GDR}}$ $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $10.7$ [6.2, 15.7]
$\textsf{GA}_{\textsf{KLR}}$ $16.0$ [14.8, 17.2] $21.9$ [16.4, 27.7] $37.2$ [29.5, 45.0]
NPO $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0]
$\textsf{NPO}_{\textsf{GDR}}$ $0.0$ [0.0, 0.0] $0.0$ [0.0, 0.0] $22.8$ [16.1, 30.1]
$\textsf{NPO}_{\textsf{KLR}}$ $17.0$ [15.7, 18.2] $25.0$ [19.0, 31.5] $44.6$ [36.5, 52.8]
Task Vector $99.7$ [99.6, 99.8] $52.4$ [45.0, 59.7] $64.7$ [57.1, 71.8]
WHP $18.0$ [16.4, 19.7] $55.7$ [48.6, 62.8] $63.6$ [56.3, 70.9]
Appendix D Dataset Details
GPT-generated QA pairs. We begin the generation by partitioning the Verbatim text of each corpus into a set of 2048-token excerpts using LLaMA-2’s tokenizer. For each QA pair to generate, we randomly sample an excerpt from this set and prompt GPT-4 (gpt-4o-2024-05-13) to create a JSON object with two fields: “question” (a question that can only be answered using specific information from the excerpt) and “answer” (an answer to the “question” extracted verbatim from the excerpt). We validate and exclude any pairs whose answers cannot be found verbatim in their corresponding excerpts. This verbatim requirement ensures that our Knowledge set is used precisely to evaluate the model’s ability to correctly associate questions with relevant portions of the training data.
For each QA pair to generate, we initiate a new conversation with GPT-4 with its corresponding excerpt. The instruction begins with a system prompt that specifies the desired format of generated QA pairs as follows:
System Prompt for Generating QAs with GPT-4 You will be provided with an excerpt of text. Your goal is to create a question-answer pair that assesses reading comprehension and memorization, ensuring that the question can only be answered using details from the excerpt. Please submit your response in a JSON format with the following fields: - “question”: A single question related to the excerpt. The question should be specific enough that it does not allow for an answer other than the one you provide. In particular, it should not be answerable based on common knowledge alone. Also, a few words extracted from the excerpt must suffice in answering this question. - “answer”: A precise answer extracted verbatim, character-by-character from the excerpt. The answer to this question must be short, phrase-level at most. The length of the extraction should be minimal, providing the smallest span of the excerpt that completely and efficiently answers the question.
We then present the excerpt as a user prompt to the model and collect the generated QA pairs. Here are two example generated QA pairs from the Knowledge set of News:
QA Pair Generated by GPT-4: Example #1 Excerpt (User prompt): …According to the Stockholm International Peace Research Institute (SIPRI), the US accounted for 69% of Israel’s arms imports between 2019 and 2023… Question: According to the Stockholm International Peace Research Institute (SIPRI), what percentage of Israel’s arms imports between 2019 and 2023 came from the US? Answer: 69%
QA Pair Generated by GPT-4: Example #2 Excerpt (User prompt): …Wednesday’s event will be moderated by tech entrepreneur David Sacks, a close ally of the Tesla founder and a supporter of Mr DeSantis… Question: Who will moderate Wednesday’s Twitter Spaces event featuring Mr DeSantis? Answer: tech entrepreneur David Sacks
Dataset segmentation. Appendix D shows examples from MUSE and Table 8 presents detailed statistics for MUSE. For both the News and Books datasets, we include the type of documents along with the number of tokens in each dataset. Additionally, MUSE incorporates $\mathcal{D}_{\textrm{retain}}^{(\text{reg})}$, a distinct retain set which is seen by $f_{\textrm{target}}$ but not included in $\mathcal{D}_{\textrm{forget}}$. This set is used exclusively with the GDR and KLR regularizers discussed. To ensure that regularized methods do not directly optimize towards the evaluation set $\mathcal{D}_{\textrm{retain}}$ , $\mathcal{D}_{\textrm{retain}}^{(\text{reg})}$ is kept disjoint from $\mathcal{D}_{\textrm{retain}}$.
Table 7: Examples of MUSE. Each corpus has Verbatim text and Knowledge sets (QA pairs derived from the original text) for evaluating verbatim and knowledge memorization. In News, $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{retain}}$ are two disjoint sets of news articles. In Books, $\mathcal{D}_{\textrm{forget}}$ is the Harry Potter book series while $\mathcal{D}_{\textrm{retain}}$ consists of wiki articles about the series. The sizes of the forget and retain sets are reported in tokens in ().
Corpus Forget Set Retain Set
News Article (0.8 M tokens) News Article (1.6 M tokens)
[…]
A: A father whose 12-year-old son
Harry Potter Books (1.1 M tokens) Harry Potter FanWiki (0.5 M tokens)
[…]
Portuguese for ‘open’.
[…]
A: Portuguese
[…]
News News Articles (3.3M) News Articles (1.6M) News Articles (1.6M) News Articles (2.0M)
Books Harry Potter Books (1.1M) Harry Potter FanWiki (0.5M) Harry Potter FanWiki (0.2M) Harry Potter Books (0.6M)